# Probabilistic Predictions of Process-Induced Deformation in Carbon/Epoxy Composites Using a Deep Operator Network

### v1

@@ 7 Optimization of the cure temperature profile
Let $DoC(t;T)$ denote the degree of cure and $u(t;T)$ the magnitude of process-induced deformation predicted by the trained DeepONet surrogate under a given temperature schedule $T(\cdot)$. The optimization problem is
@@ 8 Summary
The validated model is then used to generate a diverse dataset of PID responses over a broad family of non-isothermal cure cycles. Using this dataset, we train a FiLM-enhanced DeepONet to predict the time histories of degree of cure, viscosity, and deformation for arbitrary temperature profiles, with Feature-wise Linear Modulation allowing the branch network to account for external parameters such as the initial degree of cure. To incorporate sparse experimental information, we apply transfer learning in which only the final layer of the pretrained network is updated using measured terminal deformation, while the underlying operator learned from simulation data is kept fixed.

### v2

@@ 7 Optimization of the cure temperature profile
Let $DoC(t;T)$ denote the DoC and $u(t;T)$ the magnitude of process-induced deformation predicted by the trained DeepONet surrogate under a given temperature schedule $T(\cdot)$. The optimization problem is
@@ 8 Summary
The validated model is then used to generate a diverse dataset of PID responses over a broad family of non-isothermal cure cycles. Using this dataset, we train a FiLM-enhanced DeepONet to predict the time histories of DoC, viscosity, and deformation for arbitrary temperature profiles, with Feature-wise Linear Modulation allowing the branch network to account for external parameters such as the initial DoC. To incorporate sparse experimental information, we apply transfer learning in which only the final layer of the pretrained network is updated using measured terminal deformation, while the underlying operator learned from simulation data is kept fixed.